OpenStudy (anonymous):

Simplify the fraction to lowest terms: During one evening, there were 12 sitcoms on television and 6 movies. What fractional part of the shows were sitcoms? What fractional part were movies

OpenStudy (anonymous):

there is a total of 18 12/18= 2/3 were sitcoms 6/18= 1/3 were movies